OpenStudy (anonymous):

Solve the system y = x + 2 and y = -x - 3 by using graph paper or graphing technology.

OpenStudy (anonymous):

That site is very helpful, http://www.wolframalpha.com/

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Add the two together. y=x+2 y=-x-3 2y=-1 y=-1/2 -1/2=x+2 -5/2=x

OpenStudy (anonymous):

i need the points.. here are the options (2 , 4) (-1 , -2) (-2.5 , -0.5) (3.5 , -4.5)

OpenStudy (anonymous):

The third one. (-5/2 =-2.5) (-1/2=0.5).

OpenStudy (anonymous):

thank you !

OpenStudy (anonymous):

^(-0.5)

OpenStudy (anonymous):

No problem.